how do you find the fill in of a rectangle if the length is 5(1/4) and the width is 9(2/5)
Answer by checkley77(12844)

Don't know what the "fill" is?
5.25*9.4=49.35 is the [area] of the rectangle.
2*5.25+2*9.4=10.5+18.8=29.3 is the [perimeter].
5.25^2+9.4^2=x^2
27.5625+88.36=x^2
115.9225=x^2
x=sqrt115.9225
x=10.7667 is the [diagonal].